OpenStudy (anonymous):

A 4.0 cm diameter crankshaft turns at 2400 revolutions per minute.What is the speed of a point on the surface of the crankshaft?

OpenStudy (anonymous):

(4*400/3600)*2*pi

OpenStudy (anonymous):

solve it and u get the answer in m/sec

OpenStudy (anonymous):

d = 2r d = 4.0 cm = 4.0x10^-2 m r = d/2 = (4.0x10^-2 m)/2 = 2.0x10^-2m rpm=1/60 rev/s 2400 rpm =2400/60 rev/s = 40 rev/s \[1 rev = 2\pi rad\]\[40 rev/s = 80 \pi rad/s\] \[s=r \theta\]\[ds/dt=r (d \theta/dt)\]\[v=r \omega\]\[v = (2.0\times10^{-2} m) \times (80\pi rev/s) = 5.0 m/s\]